A block and tackle system of 5 pulleys is capable of raising load of 300N vertically upward when an effort of 75 N is applied to it downward. Calculate the efficiency of the system
1
Expert's answer
2019-02-14T10:28:37-0500

Calculate the mechanical advantage: it equals load over effort:


"MA=L\/E=300\/75=4."


Calculate the velocity ratio: it equals to the number of pulleys:


"VR=N_\text{p}=5."


Calculate the efficiency:


"\eta=\frac{MA}{VR}\cdot 100\% = \frac{4}{5}\cdot 100\% =80 \%."